𝒄(𝒙) = π₯𝐧 𝒙

πŸ‘ βˆ’ πŸ’


1
Expert's answer
2021-05-07T09:45:19-0400

Ans:-

"c(x)=lnx^3-4"

"\Rightarrow c(x)= 3\times lnx -4"

We know that "lnx=(x-1)-\dfrac{(x-1)^2}{2}+\dfrac{(x-1)^3}{3}-\dfrac{(x-1)^4}{4}+........"

So

"\Rightarrow" "c(x)=3\times[(x-1)-\dfrac{(x-1)^2}{2}+\dfrac{(x-1)^3}{3}-\dfrac{(x-1)^4}{4}+........]-4"